OpenStudy (anonymous):

Simplify. Show your work. 5 1/3 +(-3 9/18)

OpenStudy (shamim):

\[\frac{ 51 }{ 3 }+(-\frac{ 39 }{ 18 })=\frac{ 51 }{ 3 }-\frac{ 39 }{ 18 }=\frac{ 51 }{ 3 }-\frac{ 13 }{ 6}\]

OpenStudy (anonymous):